A 0.882 g sample of magnesium chloride dissolves in 91 g of water in a flask. Mg2+ What is the mol of CI? Enter to 4 decimal places.
Answer #1


Molar mass of MgCl2,
MM = 1*MM(Mg) + 2*MM(Cl)
= 1*24.31 + 2*35.45
= 95.21 g/mol


mass(MgCl2)= 0.882 g

use:
number of mol of MgCl2,
n = mass of MgCl2/molar mass of MgCl2
=(0.882 g)/(95.21 g/mol)
= 9.264*10^-3 mol

1 mol of MgCl2 gives 2 moles of Cl-
So,
moles of Cl- = 2*number of mol of MgCl2
= 2*9.264*10^-3 mol
= 1.85*10^-2 mol
Answer: 1.85*10^-2 mol